OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN! Match each equation with its solution 1.8(x + 5) = 48 A) x = 9 2.5(2x – 3) = 5 B) x = 3 3.4(x – 8) = 4 C) x = 1 4.3(3x + 1) = 30 D) x = 2

OpenStudy (anonymous):

1. 8(x + 5) = 48 would be C) x = 1

OpenStudy (anonymous):

2. 5(2x - 3) = 5 would be D. x = 2

OpenStudy (anonymous):

3. 4(x - 8) = 4 would be A) x = 9

OpenStudy (anonymous):

4. 3(3x + 1) = 30 would be B x = 3. if you want step by step just ask

OpenStudy (anonymous):

Thanks!